Public opinion reported that 7 % of Americans are afraid of being alone in the house at night .If random examples of 15 American selected .Find the probability that exactly 4 Americans are afraid of being alone.
Answer #1

Here, n = 15, p = 0.07, (1 - p) = 0.93 and x = 4
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 4)
P(X = 4) = 15C4 * 0.07^4 * 0.93^11
P(X = 4) = 0.0148